Operations Report – January 5th

Operations Report

Crew 219 Operations Report 05-01-2020

Sol: 0

Name of person filing report: Alejandro Luis Perez

Non-nominal systems: N/A

Generator : Run

Hours run : 13

From what time last night: 18:00

To what time this morning: 07:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before generator is run at night): 67%

Diesel Reading – 73%

Station Propane Reading- 57%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 55

Water Meter (units)- 0147363,4

Water (static Tank) (gal) – 550

Static to loft Pump used – Yes

Water in Green Hab (gal): 150 and still filling before the beginning
of the mission

Water in Science Dome (gal): 0

Toilet Tank Emptied – Yes

Diemos rover used: Still in the Workshop


Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director


Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

Hours: 122.9

Beginning Charge: 100%

End Charge: 65%

Currently Charging: Yes

Opportunity Rover used: Still in the workshop


Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Hours: 128.3

Beginning Charge: 100%

Ending Charge: 70%

Currently Charging: Yes

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use: N/A

Oil Added? No

# hours the ATV’s were used today: 0

Notes on ATVs: N/A

Habcar used and why, where? Yes, to fill static tank and Greenhab tank
from hollow mountain Gas station

CrewCar used and why, where? Yes, to bring Crew 218 back to grand
junction. Currently in Grand junction.

General Notes and Comments: installation of the lower deck heater, but
that it won’t be operational until positioned on the right location

Summary of the internet: Not Responsive

Summary of Suit and Radios: Radio # 8 was not receiving signal. There
are currently 2 suit helmets that need repair. One has a broken head
pad panel that does not stay in place. The other is missing the
right-side connection piece for air.

Summary of Hab: The pipe connecting static tank to loft tank has to be
defrosted by using a combination of Hairdryer to the outside pipe,
warm water poured into the feed line, and re-taping the insulation
around the outside pipe.

Summary of Science Dome operations: Adjusted latching mechanism on
Science Dome door. The door now latches closed.

Summary of Ram operations: Nominal

Summary of any observatory issues: Previous Crew had replaced both the
batteries and a new Fire alarm unit. However, the alarm was
consistently going off. According to the unit the system does not
operate in conditions under 40 °F. And there is no Co alarm in there
as well.

Summary of Health and safety issues: Smoke and Co alarm have been
replaced in Greenhab. All ABC fire extinguisher are out of date, the
observatory does not have a fire extinguisher, BC fire extinguisher in
RAM has no date and is broken and three good aerosol fire

Questions, concerns, and requests to Mission Support: Light bulbs
missing in 4 crew rooms, Max watts is 13.

Copyright © The Mars Society. All rights reserved. | Main Site